OpenStudy (anonymous):

HELP ASAP, PLEASE RIGHT NOW. CONIC SECTIONS, GOING TO SCHOOL IN A FEW MINUTES. 1. For 9y^2-4x^2-18y+24x-63=0 (y-1)^2/4-(x-3)^2/9=1 What is the domain and range? 2. For y^2=8x, the vertex is (0,0, and focus is (2,0) and the directrix is x=-2 right? PLEASE ANSWER, I WOULD HAVE ASKED LAST NIGHT BUT OPEN STUDY WAS CLOSED! THANK YOU!

OpenStudy (e.mccormick):

2) That is correct.

OpenStudy (e.mccormick):

1) That is a hyperbola, which is not a function, so I am not sure how you would get a domain and range... However, I can tell you it has no y values between 3 and -3. There it does not exist. The x values from \(\pm \infty\) are valid, but they produce multiple y values....

OpenStudy (e.mccormick):

Oops.. wait a sec, 3 and -1 on the y. Still not awake yet.

OpenStudy (anonymous):

Thank you so much for helping. I'm not sure how to find the domain and range either but I think it has something to do with the x-values or vertices.

OpenStudy (e.mccormick):

It has vertices at (3,3) and (3,-1) with nothing between them being valid. So probaly they mean a domain of \((-\infty,\infty\)) and a range of \((-\infty,-1)(3,\infty)\)

OpenStudy (anonymous):

Thank you so much! You have no idea how much you've saved me! :D

OpenStudy (e.mccormick):

Actually, inclusive. so -1]U[3

OpenStudy (e.mccormick):

Because the -1 and 3 are on the curve.

OpenStudy (e.mccormick):

\((-\infty,-1]\cup[3,\infty)\) Yah, that would be the proper notation.